OpenStudy (anonymous):

prove only using the definition that lim as x approaches 0 of (x+1)^3 = 1. So I have "Let epsilon>0. Then choose some delta>0(need to find) such that if 0

OpenStudy (anonymous):

Can you please help me?: http://openstudy.com/updates/526082dae4b0c78c62eade37

OpenStudy (bahrom7893):

I have notes somewhere on how to do this...

OpenStudy (bahrom7893):

Give me a few minutes..

OpenStudy (anonymous):

that would help so much, I think that it is just a lot of algebra manipulation and triangle identity but not sure how to manipulate it

OpenStudy (bahrom7893):

You're right.. it's been a while since i did this.

OpenStudy (bahrom7893):

Ok, here goes.. I'm gonna use e for epsilon and d for delta.

OpenStudy (bahrom7893):

Given e>0, find d>0, such that when 0<|x-x0|<d, |f(x) - L|<e \[\lim_{x \rightarrow 0}(x+1)^3=1\]

OpenStudy (bahrom7893):

Taking cube roots on both sides, we end up with:\[\lim_{x \rightarrow 0}(x+1)=1\]

OpenStudy (bahrom7893):

So now we know that:\[x_0=-1, \space L=1\]

OpenStudy (anonymous):

I do know that I have to start with l(x+1)^3-1l<e then manipulate to somehow get lxl<something*e

OpenStudy (bahrom7893):

Yea.. im just finding this hard to follow while typing.. Hang on I'll attach my notes, look at page 4, i have like 3-4 examples

OpenStudy (anonymous):

thanks for your time but I guess i just need to keep trying to rearrange it until I get what I need

OpenStudy (anonymous):

Hey did you get it?

OpenStudy (anonymous):

nope still working on it

OpenStudy (anonymous):

Don't write all the symbols stick to the stuff that is important \[\begin{split} |x-0|&<\delta &\implies |(x+1)^3-1|&<\epsilon \\ |x|&<\delta &\implies |(x^3+3x^2+3x+1)-1|&<\epsilon \\ |x|&<\delta &\implies |x^3+3x^2+3x|&<\epsilon \\ |x|&<\delta &\implies |x(x^2+3x+3)|&<\epsilon \\ |x|&<\delta &\implies |x||x^2+3x+3|&<\epsilon \\ \end{split} \]

OpenStudy (anonymous):

yep got that far then I don't know what to do

OpenStudy (anonymous):

Pick a maximum delta range you want to allow. Usually \(1\) is decent. Then figure out what \(|x^2+3x+3|\) ranges for that value.

OpenStudy (bahrom7893):

I think at this point you pick an arbitrary number for delta

OpenStudy (bahrom7893):

lol wio got there before me..

OpenStudy (anonymous):

so I would say let -1<x<1, then plug in -1 and 1 into the quadratic

OpenStudy (anonymous):

Yeah, you want to plug in -1 and 1, and any other number you might think would be the maximum.

OpenStudy (anonymous):

its probably 1 since we are approaching 0

OpenStudy (anonymous):

It looks like the highest it will be is when \(x\) is \(1\).

OpenStudy (anonymous):

It's an upwards pending parabola, so the further from it's vertex you get the higher it is.

OpenStudy (anonymous):

does that really matter since we are looking at values arbitrarily close to 0

OpenStudy (anonymous):

No, I'm giving you rigorous proof that \(x=1\) is the absolute maximum on the inverval \(-1<x<1\)

OpenStudy (anonymous):

In which case \[ |(1)^2+3(1)+3| = 7 \]

OpenStudy (anonymous):

oh ok. then we can say to choose d=min?(1,e/7)

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Well, we'd first say \[ 7|x|<\epsilon\\ |x|<\frac \epsilon 7 \]

OpenStudy (anonymous):

You want to be thorough on these, it will help you understand what you are doing better.

OpenStudy (anonymous):

right, ok thank you so much. I am never too good when choosing mins and maxs

OpenStudy (anonymous):

It is also very easy to make an error in logic and then "prove" something that is false.

OpenStudy (anonymous):

So being thorough keeps your assumptions in check. Anyway good luck.